What is the Bond Order of Li(2)(minus) anion molecule?


please explain how???

Grade:upto college level

6 Answers

Akash Kumar Dutta
10 years ago

.5

mohit yadav
10 years ago

 its bond order is o.5

pushpendra bansal
10 years ago

0.5

in molecular orbital theory bond order =1/2 (electron in b.m.o. - electron in ab.m.o)

 

kamal chandra
10 years ago

 The bond order formed by LI 2 is BO = ½ (4 – 2) = 1 Li–Li single bond. A molecul...

KUMAR PRASSU
10 years ago

bond order =1/2(Nb-Na) where Nb= number of bonding orbitals=4

                                            Na=number of non bonding orbitals=2

       B.O=0.5

hope you like it plss approve

FITJEE
10 years ago

No. of bonding elecrtons =3 (valence), anti bonding =0 (given its state)

thus =bond order =3/2=1.2